OpenStudy (anonymous):

WILL GIVE MEDAL!!! At their going-out-of-business sale, Bill's Bikes sold 34 bikes and had 50 left. Which equation could be used to find x, the number of bikes the store had originally?

OpenStudy (uri):

@lordoftanks Where are the equations?

OpenStudy (anonymous):

a x + 50 = -34 b x + 34 = 50 c x - 34 = -50 d x - 50 = 34

OpenStudy (uri):

x=34+50 duh.

OpenStudy (anonymous):

u want medal?

OpenStudy (anonymous):

lol of course you do :)

OpenStudy (uri):

or option d.

OpenStudy (uri):

same shiz

OpenStudy (anonymous):

thanx